OpenStudy (anonymous):

evaluate each function at x=5/3. f(x)=x^2+3-2 explain

OpenStudy (anonymous):

i am guessing there is a typo here and it is really \[f(x)=x^2+3x-2\]

OpenStudy (anonymous):

yes thats it

hero (hero):

In other words, evaluate \(f\left(\frac{5}{3}\right)\)

OpenStudy (anonymous):

yes

hero (hero):

Okay so \(f\left(\dfrac{5}{3}\right) = \left(\dfrac{5}{3}\right)^2 + 3\left(\dfrac{5}{3}\right) -2 \)

hero (hero):

Can you simplify: \(\left(\dfrac{5}{3}\right)^2 + 3\left(\dfrac{5}{3}\right) -2\)

OpenStudy (anonymous):

nope